Ques:2
Use Differentiation of Vector-Valued Functions to calculate the derivative of each of the
following functions.
a. r(t)=(6t+8)i+(4t2+2t−3)jr(t)=(6t+8)i+(4t2+2t−3)j
b. r(t)=3costi+4sintj
1
Expert's answer
2021-07-11T17:22:26-0400

Let us use Differentiation of Vector-Valued Functions to calculate the derivative of each of the

following functions.


a. "r(t)=(6t+8)i+(4t^2+2t\u22123)j"


It follows that "r'(t)=6i+(8t+2) j"


b. "r(t)=3\cos t i+4\sin t j"


It follows that "r'(t)=-3\sin ti+4\cos t j"